A sprinter starts from rest, and reaches their top speed of 12 m/s in 3 seconds.
Part 1) What is the acceleration of the sprinter?
Part 2) What is the average velocity of the sprinter during those 3 seconds?
Part 3) How far did the sprinter travel during those 3 seconds?
1
Expert's answer
2020-10-16T10:57:52-0400

1)


"a=\frac{v}{t}=\frac{12}{3}=4\frac{m}{s^2}"

2)


"V_{av}=0.5(v_i+v_f)=0.5(0+12)=6\frac{m}{s}"

3)


"s=V_{av}t=(6)(3)=18\ m"